What is the pH for the solution made from 11.10 mL of 0.278 M HCl solution mixed with 70.90 mL of 0.1300 M ammonia solution? Note that the pKa for NH4+ is 9.241

1)9.24

2)8.77

3)9.54

4)9.72

5)8.94



1
Expert's answer
2021-06-07T03:24:23-0400

"PH= pka + log( base\/ acid)"


PH=9.241+ log(0.130/0.278)

= 9.241- 0.33

pH= 8.94